Question 18 Assume TC = 363 +350 +3Q2. At what Q is average cost at its minimum? Enter as a value. < Previous Not sa
Answer #1

At minimum. da ta dTC da =0 d (363 +352 +3Q2) =0 dQ 35 + 60 =0 * Q=1-5.83]